How many rows appear in a truth table for each of these

compound propositions?

a) (q -+ -'p) v (-'P -+ -.q)

b) (p v -.t) /\ (p v -'s)

c) (p -+ r) V (-,s -+ -.t) v (-,u -+ v)

d) (p /\ r /\ s) V (q /\ t) V (r /\ -.t)


1
Expert's answer
2022-03-13T18:37:33-0400

a. "2^2=4"

b. "2^{3}=8"

c. "2^6=64"

d. "2^5=32"